OpenStudy (anonymous):

3+2/3y=1/2x in standard form

jigglypuff314 (jigglypuff314):

standard form for a line is ax + by = c you want the values with variables all on one side so subtract both sides by (1/2)x then subtract both sides by 3

OpenStudy (anonymous):

so is that just 2/3y+-1/2x+3?

OpenStudy (shamil98):

you want whole numbers in standard form @jigglypuff314 not fractions

OpenStudy (anonymous):

never?

OpenStudy (shamil98):

-1/2 x + 2/3y = -3 multiply the equation by a value to get them as whole numbers.

OpenStudy (anonymous):

so what is the answer?

OpenStudy (shamil98):

what value makes the fractions whole numbers?

OpenStudy (shamil98):

what value multiplied*

OpenStudy (anonymous):

6?

jigglypuff314 (jigglypuff314):

yes so you would get -3x + 4y = -18

OpenStudy (anonymous):

does the y not have to be by itself?

OpenStudy (shamil98):

yes 6 is the value you multiply by it, standard form is ax + by = c so no you dont get y byitself that is slope-intercept form which is y = mx +b